OpenStudy (anonymous):

Once upon a time there was an old man who owned a number of camels. He was considering giving the camels to members of his family and thought: “Perhaps I should let my two sons have the camels – no, there would be one camel left over. Perhaps I should divide the camels among all five of my children – no, there would be two camels left over. Perhaps I should share the camels among my 17 children and grandchildren – no, there would be 3 camels left over. I had better just keep the camels all for myself!” How many camels does the old man own?

OpenStudy (anonymous):

would it be 23 assuming that each person in the last scenario gets 1 camel plus the 6 remaining camels? 17+6=23?

OpenStudy (anonymous):

Let the number of camels be x, when divided by 17, it leaves a remainder of 3, so therefore x=17n+3, where n is a natural number Also since the there is still one camel left when x is divided among two, it must be an ODD number Thus the 'n' would be an even number, because odd into even will give u even and +3 will make it odd so your possibilities can be n=2,4,6,8,....

OpenStudy (anonymous):

At this point you can just manually check...

OpenStudy (anonymous):

so there isn't just one definite answer?

OpenStudy (anonymous):

Nope, there is, try one of the even values of n and see if it fits with the division among 5 people

OpenStudy (anonymous):

so would I do 17*even number divided by 5?

OpenStudy (anonymous):

Try the even values, whatever answer you get, divide it by 5, if u get the remainder as 2 then you have the answer

OpenStudy (anonymous):

oh okay! so like 17*2+3/5?

OpenStudy (anonymous):

For example \[17 \times 6 +3=105\]\[105/5=21\] remainder is not 2 in this case

OpenStudy (anonymous):

oh i see! doing the math now!!

OpenStudy (anonymous):

I havent gotten anything with R2 :o

OpenStudy (anonymous):

17*2+3=37 37/5=7.4

OpenStudy (anonymous):

and 2 is the smallest even number

OpenStudy (anonymous):

How much is the remainder when 37 is divided by 5?You're telling me the quotient..(7.4)

OpenStudy (anonymous):

OH DUH

OpenStudy (anonymous):

I'm sorry I'm so slow, i was using a calculator instead of doing the math

OpenStudy (anonymous):

its 7 R2

OpenStudy (anonymous):

Yep, so 37 when divided among 2, everyone gets 18 camels while 1 is left over, 37 when divided among 5, everyone gets 7 camels while 2 are left over, 17 when divided among 17, everyone gets 2 camels while 3 are left over, Is everything satisfied as per the question?See for yourself

OpenStudy (anonymous):

wow that is so cool how we were able to get the number of camels just by that information! i felt like it was impossible! thank you so so much!! and thanks for being so patient :)

OpenStudy (anonymous):

:)